How many grams are in 1.48 × 107 μg?
Answer #1

conversion factor:

1 ug = 1.0*10^-6 g

So,

1.48*10^7 ug = 1.48*10^7 * 1.0*10^-6 g

= 14.8 g

Answer: 14.8 g